5. Juli 2004. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), EB = Erziehungswiss. Bibl., FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB = Informatikbibliothek, IB Nr. = Institutsbibliothek Nr., LB = HB– Lehrbuchsammlung, LS = HB–Lesesaal

Literatur zu Henselmethoden (Rechnen via Homomorphismen)

[1] A. G. Akritas. Elements of Computer Algebra with Applications. Wiley, 1989. MB: 14888. S. 320–327.

[2] E. J. Barbeau. Polynomials. Problem Books in Mathematics. Springer–Verlag, 1989. MB: 17 762. S. 102–104 und 172–173.

[3] K. Hensel. Eine neue Theorie der algebraischen Zahlen. Math. Zeitschr., 2:433–452, 1918.

[4] N. Koblitz. p–adic Numbers, p–adic Analysis, and Zeta–Functions. GTM 58. Springer–Verlag, 1977. HB: 58 Bb 1244, MB: 9544. S. 16–18.

[5] H. Lüneburg. Rechnen via Homomorphismen. Bayreuther Math. Schriften, 21:279–294, 1986. MB: 15 936 n. S. 286 ff.

[6] Sibylla Priess-Crampe and Paulo Ribenboim. A general Hensel’s lemma. J. of Algebra, 232(1):269–281, 2000. Hensel’s lemma for several polynomials.

[7] Sibylla Priess-Crampe and Paulo Ribenboim. Fixed point and attractor theorems for ultrametric spaces. Forum Math., 12:53–64, 2000. MB-Link. Zbl 0944.54031. Anw. bei Hensel-Methoden; easier proffs of classical theorems.

[8] H. Zassenhaus. On Hensel factorizations I. J. Number Theory, 1:291–311, 1969. MB: Z 106.

[9] H. Zassenhaus. A remark on the Hensel factorization method. Math. of Computation, 32:287–292, 1978.